OpenStudy (anonymous):

Help please!!! Medals will be given The function below shows the number of car owners f(t), in thousand, in a city in different years t: f(t) = 0.25t2 - 0.5t + 3.5 The average rate of change of f(t) from t = 2 to t = 6 is ______ thousand per year. Numerical Answers Expected!

OpenStudy (anonymous):

19

OpenStudy (anonymous):

ok thanks! how'd you get that?

OpenStudy (anonymous):

did the work

OpenStudy (anonymous):

\[\frac{f(6)-f(2)}{6-2}\] is what you have to compute

OpenStudy (anonymous):

So is 19 right? @satellite73

OpenStudy (anonymous):

should be

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

i get \[f(6) = 9.5\] and \[f(2) = 3.5\] and so \[\frac{f(6)-f(2)}{6-2}=\frac{9.5-3.5}{4}=\frac{6}{4}=\frac{3}{2}=1.5\]

OpenStudy (anonymous):

So what do you think the correct answer is?

OpenStudy (anonymous):

no idea where the 19 came from i think the correct answer is the one i just wrote

OpenStudy (anonymous):

So it would be 1.5?

OpenStudy (anonymous):

@satellite73